OpenStudy (anonymous):

Solve: 5x − 10 + 12 = 22

OpenStudy (anonymous):

what is -10+12

OpenStudy (anonymous):

2

OpenStudy (anonymous):

okay so we have 5x +2 = 22

OpenStudy (anonymous):

what is 5x+2-2=22-2 ?

OpenStudy (anonymous):

5x=20 ohh so x=4?

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

thank you